2. [6 marks] Are the following functions O(n)? Justify your answer. a) n log n b) f(n) = Vn (log n)
Answer #1
a)
f(n) = O(g(n)) means there are positive constants c and n0, such that 0 ≤ f(n) ≤ cg(n) for all n ≥ n0

nlog(n) = O(n^2)
=>  nlog(n) <= c(n^2)
Let's assume c = 1
=>  nlog(n) <= c(n^2)
=>  nlog(n) <= 1(n^2)
=>  nlog(n) <= n^2
=>  log(n) <= n
it's true for all n >= 1

so, nlog(n) = O(n^2) given c = 1 and n0 = 1, using the definition of Big-O

b)
f(n) = O(g(n)) means there are positive constants c and n0, such that 0 ≤ f(n) ≤ cg(n) for all n ≥ n0

√n log(n) = O(n^2)
=>  √n log(n) <= c(n^2)
Let's assume c = 1
=>  √n log(n) <= c(n^2)
=>  √n log(n) <= 1(n^2)
=>  log(n) <= n√n
it's true for all n >= 1

so, √n log(n) = O(n^2) given c = 1 and n0 = 1, using the definition of Big-O